Brain volumes of 20 brains have mean of 1081.3 and standard deviation of 128.4. What are the significantly low and high values? Would a brain volume of 1358.1 be significantly high?
1
Expert's answer
2021-10-14T18:14:40-0400

"\mu=1081.3"

"\sigma=128.4"

Using the range rule of thumb, significantly low/high values are given as,

Significantly low value="\mu-2*\sigma"

Significantly low value =1081.3-(2*128.4)=824.5

Significantly high value="\mu+2*\sigma"

Significantly high value=1081.3-(2*128.4)=1338.1

The limit of values is 824.5 to 1338.1.

A value below 824.5 is considered low and a value above 1338.1 is considered high.

Therefore, the value of 1358.1is more than the upper limit, thus it is significantly high